OpenStudy (anonymous):

A box contains 12 plastic forks and 6 plastic knives. If two utensils are chosen at random from the box without replacement, what is the probability that they are different? A) 13 17 B) 4 17 C) 8 17 D) 9 17 PLEASEE EXPLAIN HOW TO DO THIS. I don't just want the answer :) Thank you VERY much.

OpenStudy (anonymous):

So there are 18 total utensils in the box. What's the probability that a fork is drawn first?

OpenStudy (anonymous):

12 over 18?

OpenStudy (anonymous):

Correct. It can be reduced to 2/3. The fork is NOT replaced so there are now 17 utensils in the box. What's the probability that a knife is chosen next?

OpenStudy (anonymous):

6 over 17

OpenStudy (anonymous):

Very good. So the probability that a fork is drawn first and a knife is drawn second is the product of these two individual probabilities, i.e. 2/3 x 6/17. What do you get?

OpenStudy (anonymous):

Do i cross multiply or go straight across?

OpenStudy (anonymous):

To multiply fractions, multiply the numerators together and multiply denominators together

OpenStudy (anonymous):

12 over 14 ?

OpenStudy (anonymous):

Not quite.\[\frac{ 2 }{ 3 } \times \frac{ 6 }{ 17 } = \frac{ 2 \times 6 }{ 3 \times 17 } = ?\]And don't forget to reduce the answer to lowest terms.

OpenStudy (anonymous):

OH, duh for me, 12/21. Sorry :p

OpenStudy (anonymous):

Nope. Try again. Use a calculator.

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

What's 3 x 17 ?

OpenStudy (anonymous):

Wait i thought we were timing 7 times 3?

OpenStudy (anonymous):

Where does the 7 come from? The denominator is 3 x 17

OpenStudy (anonymous):

i AM SO SORRY. Rough morning. okay so 17 times 3 is 51 so its 12/51?

OpenStudy (anonymous):

Very good. But this can be reduced. 3 is a common factor, so divide both numerator and denominator by 3. What do you get?

OpenStudy (anonymous):

4/17?

OpenStudy (anonymous):

Excellent. So the probability of choosing a fork first and then a knife is 4/17. But we're not done yet. Do you know why?

OpenStudy (anonymous):

Ummm, well I dont think it can be simplified further... So i'm not really sure. Because it wants to know the probability that they are different??(Again sorry, rough morning an im horrible at math)

OpenStudy (anonymous):

No problem. You're right that it can't be simplified further. But, there is another possibility that satisfies the question. The question asks for the probability that the two choices will be different, and you calculated the probability of choosing a fork, then a knife. But what about choosing a knife first, then a fork. You also need to calculate that probability. Understand?

OpenStudy (anonymous):

Yes, but i thought it would just switch the fractions that we multiplied around and i would get the same answer

OpenStudy (anonymous):

Maybe, but let's go through it.

OpenStudy (anonymous):

So, what's the probability of choosing a knife out of the box first?

OpenStudy (anonymous):

6/17

OpenStudy (anonymous):

Remember, when you begin, there are 18 utensils in the box. Try again.

OpenStudy (anonymous):

6/18 which simplifies to 1/3

OpenStudy (anonymous):

Exactly. Now, there are 17 utensils in the box. What's the probability of choosing a fork second?

OpenStudy (anonymous):

12/17

OpenStudy (anonymous):

Correct. Now, the probability of choosing knife then fork is the product of these two probabilities. What do you get?

OpenStudy (anonymous):

12/51

OpenStudy (anonymous):

Excellent. Now, 3 is also a common factor so this fraction can be reduced, to what?

OpenStudy (anonymous):

6/17

OpenStudy (anonymous):

Not quite. What's 12 divided by 3?

OpenStudy (anonymous):

oh, 4 sorry

OpenStudy (anonymous):

Right. So now you have everything you need. Notice that the probabilities for the two possibilities ended up the same but the individual fractions were not simply reversed. So you can't assume things, you must go through and do the calculations.

OpenStudy (anonymous):

Alright, now to answer the question. You've determined that the probability of choosing a fork first and a knife second is 4/17. AND, you've determined that the probability of choosing a knife first and a fork second is 4/17. Thses are the only two scenarios that satisfy the question. So the final answer is the SUM of these two probabilities. What is it?

OpenStudy (anonymous):

So I got 8/34 but i feel like thats wrong

OpenStudy (anonymous):

Sorry. That's not it. Remember, we're not multiplying fractions here, we're adding them. When adding fractions with a common denominator, keep the denominator the same and add the numerators. For example,\[\frac{ 1 }{ 5 } + \frac{ 3 }{ 5 } = \frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

2/6 which simplifies to 1/3???

OpenStudy (anonymous):

Nope. I'll write it out.\[\frac{ 4 }{ 17 } + \frac{ 4 }{ 17 } = ??\]

OpenStudy (anonymous):

Keep the denominator the same and add the numerators.

OpenStudy (anonymous):

8/17

OpenStudy (anonymous):

Bingo! That's your answer. Good job.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You're welcome.